# AI as Coordination-Compressing Capital:

### v1

@@ 5.4 Limitations
This paper has several limitations. First, the production function is stylized—a two-layer firm with Cobb–Douglas structure does not capture the complexity of real organizational design. In particular, the fixed factor shares ($\alpha$, $1-\alpha$) mechanically link the manager–worker gap to team size; under a CES specification with elasticity of substitution $\sigma\neq 1$, the managerial share could shrink as spans expand (if management and effective labor are substitutes), potentially attenuating or reversing the gap result. The regime fork’s qualitative ordering is robust to $\alpha$ (Table 4), but testing robustness to $\sigma$ is an important extension. Second, the simulations are illustrative: they demonstrate that the model generates the regime fork, but calibration to specific industries or firms requires richer data than is currently available. Third, the simulation models a single multi-manager firm, not a general-equilibrium economy—inter-firm competition, labor market clearing, and endogenous entry are abstracted away. Fourth, worker skills are exogenous and fixed; in practice, AI may alter the returns to skill acquisition, endogenizing the skill distribution. Fifth, the model abstracts from firm-boundary effects (Coase, 1937; Williamson, 1979)—coordination compression may shift the make-or-buy margin in ways not captured here. Sixth, coordination cost is strictly decreasing in $K_{A}$ with no floor; in practice, verification and auditing overhead from managing large numbers of concurrent processes may impose a lower bound on effective coordination friction, attenuating span expansion at high $K_{A}$ without altering the regime fork’s qualitative structure.
@@ 6 Conclusion
This paper extends the task-based framework for AI and labor by introducing coordination compression as a distinct channel through which agentic AI reshapes organizations. Agent capital ($K_{A}$) reduces the friction of managing workers, expanding spans of control, compressing hierarchies, and—under positive task creation elasticity—expanding the frontier of feasible work. The model generates five propositions and a regime fork: the same technology produces broad-based gains or superstar concentration depending on a single parameter ($\beta$) that indexes who benefits from coordination compression. Numerical simulations with heterogeneous managers and workers confirm sharp regime divergence across a $2\times 2$ parameter space. The regime fork holds economy-wide: coordination compression universally reduces unemployment and widens the manager–worker gap, but the pace of inequality reduction depends critically on $\beta$.

### v3

@@ 5.4 Limitations
This paper has several limitations. First, the production function is stylized—a two-layer firm with Cobb–Douglas structure does not capture the complexity of real organizational design. In particular, the fixed factor shares ($\alpha$, $1-\alpha$) mechanically link the manager–worker gap to team size; under a CES specification with elasticity of substitution $\sigma\neq 1$, the managerial income share could shrink as spans expand (if management and effective labor are substitutes, $\sigma>1$), potentially attenuating the gap result. Full reversal would likely require additional changes beyond the production function alone (e.g., endogenous management intensity, alternative surplus-sharing, or non-binding capacity constraints). The regime fork’s qualitative ordering is robust to $\alpha$ (Table 4), but testing robustness to $\sigma$ is an important extension. Second, the simulations are illustrative: they demonstrate that the model generates the regime fork, but calibration to specific industries or firms requires richer data than is currently available. Third, the simulation models a single multi-manager firm, not a general-equilibrium economy—inter-firm competition, labor market clearing, and endogenous entry are abstracted away. Fourth, worker skills are exogenous and fixed; in practice, AI may alter the returns to skill acquisition, endogenizing the skill distribution. Fifth, the model abstracts from firm-boundary effects (Coase, 1937; Williamson, 1979)—coordination compression may shift the make-or-buy margin in ways not captured here. Sixth, coordination cost is strictly decreasing in $K_{A}$ with no floor; in practice, verification and auditing overhead from managing large numbers of concurrent processes may impose a lower bound on effective coordination friction, attenuating span expansion at high $K_{A}$ without altering the regime fork’s qualitative structure. Seventh, the elite complementarity parameter $\beta$ performs double duty: it simultaneously governs how pre-existing skill differences translate into coordination capacity and how AI access is distributed. In particular, the formulation $K_{\text{eff},i}=K_{A}\cdot s_{i}^{\beta}$ implies that at $K_{A}=0$, all managers have identical coordination costs regardless of skill, producing a Gini of zero as a baseline normalization rather than reflecting realistic pre-existing inequality. Separating baseline managerial heterogeneity in coordination capacity from the AI complementarity rate—for instance, via $c_{i}(K_{A})=c_{0}\cdot s_{i}^{-\eta}/(1+\gamma K_{A}s_{i}^{\beta})$ where $\eta>0$ captures pre-AI skill sorting—would allow the model to distinguish between pre-existing organizational inequality and AI-induced amplification. We leave this extension for future work.
@@ 6 Conclusion
This paper extends the task-based framework for AI and labor by introducing coordination compression as a distinct channel through which agentic AI reshapes organizations. Agent capital ($K_{A}$) reduces the friction of managing workers, expanding spans of control, compressing hierarchies, and—under positive task creation elasticity—expanding the frontier of feasible work. The model generates five propositions and a regime fork: the same technology produces broad-based gains or superstar concentration depending on a single parameter ($\beta$) that indexes who benefits from coordination compression. Numerical simulations with heterogeneous managers and workers confirm sharp regime divergence across a $2\times 2$ parameter space. Within the simulated firm, the regime fork extends beyond the managerial layer: coordination compression reduces the share of unassigned workers and widens the manager–worker gap across all regimes, but the pace of inequality reduction depends critically on $\beta$. Whether these within-firm patterns aggregate to economy-wide effects depends on general-equilibrium forces—inter-firm competition, labor market clearing, and endogenous entry—that the current partial-equilibrium framework does not model.